OpenStudy (anonymous):

Solve for x. 20x - 2 = 18x - 4

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

\[20x-2=18x-4\] \[20x-18x=-4+2\] \[2x=-2\] \[x=\frac{-2}2\] \[x=-1\]

OpenStudy (anonymous):

20 x-2 = 18 x-4 Add (2-18 x) to both sides: 2 x = -2 Divide both sides by 2: x = -1